iYuko (iyuko):

Introduction to Factoring Lesson 2 Come learn with me!

iYuko (iyuko):

Factoring Trinomials in the form of \[x ^{2}+bx+c\]

iYuko (iyuko):

Suppose you wanted to factor \[x ^{2}+5x+6\]. After some trial and error you might obtain \[(x+2)(x+3)\] Or you might get discouraged and not get an answer. If you did get these factors, you could check your answer by the FOIL method.\[(x+2)(x+3)=x ^{2}+3x+2x+6\]\[=x ^{2}+5x+6\] But trial and error can be a long process There is another way.

iYuko (iyuko):

Look at the original polynomial \[x ^{2}+5x+6\] We know immediately that the answer will be in the form ( )( ). Since the first term of the polynomial is x^2, we can place an x in each parentheses, (x )(x ). Next look at the sign of the last term. The sign is positive. This means that the sign of each factor must be the same. Since the sign of the middle term is positive, we can write (x+ )(x+ ). Look at the sign of the last term. Think of factors of 6. Try 2 and 3 since 2+3=5, the coefficient of the middle term. We write (x+2)(x+3) and our factorizing is complete. We can check using the FOIL as show above.

iYuko (iyuko):

Factoring Trinomials on the Form of x^2+bx+c 1. The answer will be in (x+ )(x+ ). 2. The two numbers at the end of each set of parentheses are numbers such as: (a). When you multiply them, you get the last term, which is c. (b). When you add them, you get the coefficient of x, which is b.

iYuko (iyuko):

So far we have only factored trinomials in the form of x^2+bx+c, where b and c are positive numbers. The same procedure applies is b is negative and c is positive. Remember, is c is positive, the sign of both factors will be the same.

Nnesha (nnesha):

Nice! great wO_Ork!

Nnesha (nnesha):

if you put "when you multiply you should get the product of ac * *i guess* that would be better when leading coefficient \[\cancel{= }\]equal to one

Nnesha (nnesha):

^ huh?

OpenStudy (anonymous):

\(\huge\bf\rlap{\color{Lime}{iYuko\ is\ Awesome}}{\color{Blue}{\; iYuko\ is\ Awesome}}\)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @iYuko Lesson 1: http://openstudy.com/study#/updates/5531db15e4b05a945296afd8 \(\color{blue}{\text{End of Quote}}\) that's what i got http://prntscr.com/6v4sux lol

iYuko (iyuko):

http://openstudy.com/study#/updates/552f38fae4b01573da53b3cf Here you are.

OpenStudy (anonymous):

lel lets delete our messages and burn the evidence

OpenStudy (anonymous):

well the bad ones

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im a beginner

OpenStudy (anonymous):

WTF LOL U A PRO

OpenStudy (anonymous):

XD WTF

OpenStudy (anonymous):

U R AN IDIOT

iYuko (iyuko):

\(\bbox[5pt, white ,border:5px dotted black]{\bf\huge\color{#FF004F}{@iYuko}}\)

Nnesha (nnesha):

@Icedragon50 language .... TUTORIAL!

OpenStudy (anonymous):

@iYuko Reported.

OpenStudy (anonymous):

send it to me. i wanna learn this

iYuko (iyuko):

What? Just read it.

iYuko (iyuko):

@Preetha Can you delete the irrelevant messages?

OpenStudy (anonymous):

You sent an innocent person to a site that called them an idiot.

OpenStudy (anonymous):

Huh? Latex?

iYuko (iyuko):

\(\Huge\mathbb{\color{coral}{This~Is~LaTeX}}\)

iYuko (iyuko):

@OpiGeode

OpenStudy (anonymous):

You have a unique style of being a fail troll.